OpenStudy (anonymous):

hey! how can i find a formula for p(x) with a degree of 4, a root of multiplicity 2 at x=4 and roots of multiplicity 1 at x=0 and x= -4, also it goes through the point (5, 18)?

OpenStudy (anonymous):

If you have roots with multiplicity 2 at 4 and multiplicity 1 at x=0 and x=-4 you have something in the form of: \[ax(x-4)^2(x+4)\] The constant a SHOULD dictate the point it goes through. We have: \[18=a(5)(5-4)^2(5+4)=45a \rightarrow a=18/45\] So your polynomial is : \[\frac{18x}{45}(x-4)^2(x+4)\]